The Blue Star Manufacturing Company uses a job order costing system and normal costing. It started operations in January 2014. Total production cost for Job A14 was $35,000. Job A14 was completed, but not sold as of the end of 2014. During 2014, the total production cost (DM + DL + OH) charged to B14 was $24,000. Job B14 was still unfinished at the end of 2014. The ending balance for Materials Inventory for 2014 was $6,000. At the end of 2014, all underapplied or overapplied overhead was closed totally to Cost of Goods Sold. During 2015, Blue Start started three new jobs: A15, B15 and C15. During 2015, the company charged overhead to the product based on Direct Labor Cost. Estimated overhead for 2015 was $60,000 and estimated Direct Labor Cost for 2015 was $48,000. Direct Materials and Direct Labor Costs charged to jobs were follows: Job B14 Job A15 Job B15 Job C15 Direct Materials $10,000 $15,000 $20,000 $ 6,000 Direct Labor $12,000 $16,000 $ 8,000 $ 4,000 Direct Material purchased during 2015 cost $55,000. Actual overhead incurred during the time period was $47,000. All jobs that had been worked on during 2015 were completed at the end of the year except Job C15. Only Job A15 was still in Finished Goods Inventory as of the end of 2015. To answer the ten questions below, be sure to consider the information, as relevant, from 2014. (Don’t forget to charge overhead to jobs, where relevant, as well.)

31. Compute the Ending Materials Inventory for 2015. a. $4,000 b. $6,000 c. $8,000 d. $10,000

32. Compute the Unadjusted Cost of Goods Sold for 2015. a. $99,000 b. $75,000 c. $110,000 d. $134,000
Answer #1
Question 31
Value of ending inventory
Opening inventory 6000
Direct Material purchased during 2015 55000
Less:Direct material used
Job B14 -10000
Job A15 -15000
Job B15 -20000
Job C15 -6000
Ending inventory 10000
Therefore Option D is correct
Question 32
Cost of goods Sold
Opening inventory 6000
Add : cost of goods manufactured
Job B14=10000+12000 22000
Job A15=15000+16000 31000
Job B15=20000+8000 28000
Job C15=6000+4000 10000
Overhead 47000
Goods available for sale 144000
Less: Ending inventory 10000
Cost of goods sold 134000
Therefore Option D is correct